I need help solving literal equations for one of its variables.K=1/2mv^2 (solve for m)
Answer by Alan3354(69443)

I need help solving literal equations for one of its variables.K=1/2mv^2 (solve for m)
--------------
K=1/2mv^2 Just get m on one side of the equal sign alone.
2K = mv^2
m = 2K/v^2